### v1

# New Projection Operators for Planar Electrodynamics

Abstract
In this article, we provide a new method for obtaining the propagator of two three-dimensional models of electrodynamics (Maxwell-Lee-Wick-Chern-Simons and Maxwell-Deser-Jackiw). This method introduce a new set of projection operators. Then we perform a causality and unitarity analysis.
1  Introduction
In Quantum Field Theory (QFT) the calculation of the propagator is essential for the perturbative development in obtaining cross sections [zuber, schwartz, peskin], as well as for elucidating the nature of the excitations that arise in the interaction between particles and fields [accioly, baeta, boldo]. For this reason, it becomes important to know different methods for calculating these propagators.
It is well known that the pole structure of the propagator encodes the physical dispersion relations of the excitations, since the on-shell energies follow from the zeros of the corresponding denominators in momentum space. Likewise, the analysis of the residues at these poles-in particular, the sign of the residue in the saturated amplitude-provides a direct criterion for tree-level unitarity and is closely tied to the causal properties of the propagating modes [accioly, baeta, boldo].
In this work we present an alternative method to obtain the propagator of three-dimensional models, in particular the case of Maxwell-Lee-Wick-Chern-Simons [leewick] and Maxwell-Deser-Jackiw electrodynamics [djackiw]. The proposed method consist of the decomposition into direct sums of subspaces; this method introduce a new set of projection operators that make easier the analysis of causality and unitarity of the propagator.
We begin by presenting a general three-dimensional theory for a vectorial field, the problem in Section 2, exemplified by the case of Maxwell-Lee-Wick-Chern-Simons electrodynamics [leewick], taking as a starting point the Lagrangian and obtaining the corresponding field equations in terms of the wave operator. In the Section 3, we review some concepts and theorems of Linear Algebra, whose demonstrations are found in the Appendix. This allows us to introduce the method and present its implications in a clearer way, in addition to its application. Next, in the Section 4, we apply the method to Electrodynamics. Later, in the Section 5, we apply the method to obtain the propagator of Maxwell-Lee-Wick-Chern-Simons and Maxwell-Deser-Jackiw electrodynamics. In the Section 6, we present the causality and unitarity analysis for the models. Finally, in the Section 7, we summarize our main results and discuss possible extensions and future directions of the proposed work.
2  Problem Definition
We start with a Lagrangian in the form
$\mathcal{L}=\frac{1}{2}A^{\mu}\mathcal{O}_{\mu\nu}A^{\nu}-J_{\mu}A^{\mu},$
where the wave operator in the momentum space is
$\mathcal{O}^{\mu}_{\ \nu}(p)=\alpha(p)\,\omega^{\mu}_{\ \nu}(p)+\beta(p)\,\theta^{\mu}_{\ \nu}(p)+\gamma(p)\,S^{\mu}_{\ \nu}(p).$ (1)
Where,
$\omega^{\mu}_{\;\nu}=\frac{p^{\mu}p_{\nu}}{p^{2}},\quad\theta^{\mu}_{\;\nu}=\delta^{\mu}_{\;\nu}-\omega^{\mu}_{\;\nu}\quad\text{and}\quad S^{\mu}_{\;\nu}=\,\varepsilon^{\mu}_{\,\,\,\nu\beta}\,p^{\beta}.$
With this, the functional generator is
$Z[J]=\mathcal{N}\int\mathcal{D}A\;\exp\left\{i\int d^{3}x\mathcal{L}\right\}.$ (2)
Completing the square and performing the Gaussian integration, we obtain
$Z[J]=Z[0]\,\exp\left\{-\frac{i}{2}\int d^{3}x\,d^{3}y\;J_{\mu}(x)\,G^{\mu}{}_{\nu}(x-y)\,J^{\nu}(y)\right\}.$ (3)
The two point functions is
$\langle 0|TA_{\mu}(x)A_{\nu}(y)|0\rangle=\frac{1}{i^{2}}\frac{\delta^{2}Z[J]}{\delta J^{\mu}(x)\,\delta J^{\nu}(y)}\bigg|_{J=0}=iG_{\mu\nu}(x-y).$ (4)
Where,
$\mathcal{O}^{\alpha}{}_{\nu}\,G^{\nu}{}_{\beta}(x-y)=\delta^{\alpha}{}_{\beta}\,\delta^{(3)}(x-y).$ (5)
Therefore, the calculation of the Green’s functions is basically reduced to the inversion problem of (1), which is precisely the goal of this work. For that, we need to understand some mathematical definitions and a little more about projectors. In order to obtain the propagator, our method need some algebraic definitions that will be used for the inversion, the main ones being some results about Projectors and their relation to direct sum decomposition. The proofs of the theorems stated in this section can be found in the Appendix.
3  Mathematical Formalism
The discussion below is grounded on a general algebraic result, namely the Cayley-Hamilton theorem [lima_algebra, axler2024]. Since the vector space $V$ is finite-dimensional, any linear operator acting on it satisfies its own characteristic polynomial. When such an operator is invertible, this implies that its inverse can be written as a linear combination of the identity and its powers. In the context considered here, this theorem provides the conceptual justification for the closure properties assumed for the operators involved and explains why propagators can be consistently expressed within the same algebraic framework.
Given a linear operator $T:V\to V$, where $V$ is a finite-dimensional vector space. Suppose that $T$ is given by a linear combination:
$T=\sum_{i=1}^{n}\alpha_{i}T_{i},$ (6)
which occurs with wave operators where, as in (1), we see that $T^{-1}$, if it exists, will be a linear combination of the identity $I$ and products of the operators $T_{i}$. We are therefore motivated to determine the products $T_{i}T_{j}$.
A particularly interesting case occurs when $T_{i}T_{j}$ belong to $\text{span}\{T_{1},\ldots,T_{n}\}$ (the subspace generated by the operators $T_{i}$), that is, $T_{i}T_{j}=\sum_{k}\beta_{ijk}T_{k}$. In fact, if $T=\sum_{i}\alpha_{i}T_{i}$ is invertible, then $T^{-1}\in\text{span}\{I,T_{1},\ldots,T_{n}\}$, that is, it will be a linear combination of $I$ and of the $T_{i}$. We note that, in this case, the subspace $\text{span}\{I,T_{1},\ldots,T_{n}\}$ will be a subalgebra, that is, if $T,U\in\text{span}\{I,T_{1},\ldots,T_{n}\}$ and $T$ is invertible, then $TU,T^{-1}\in\text{span}\{I,T_{1},\ldots,T_{n}\}$.
Definition 3.1.
Let $P:V\to V$ be a linear operator. We say that $P$ is a projector if and only if $P^{2}=P$. Operators of this type are called Idempotents.
An immediate consequence is that every vector in the image $\mathrm{Im}\,P=\{Pv;v\in V\}$ of a projector $P$ is its own eigenvector of eigenvalue $1$, that is, if $v\in\mathrm{Im}\,P$, then $Pv=v$. On the other hand, every vector in the kernel $\ker P=\{v\in V;Pv=0\}$ is an eigenvector of eigenvalue 0.
Theorem 3.1.
Let $P:V\to V$ be a projector. Then one has:
$V=\mathrm{Im}\,P\oplus\ker P.$ (7)
In other words, a projector $P$ decomposes $V$ as a direct sum of its image and kernel. Moreover, $\dim(V)=\dim(\ker P)+\dim(\mathrm{Im}\,P)$.
Theorem 3.2.
Let $P$ be a projector. Then IPI-P is also a projector. Moreover,
$\ker P=\mathrm{Im}(I-P),\quad\mathrm{Im}\,P=\ker(I-P).$ (8)
Theorem 3.3.
Let $\mathcal{C}=\{P_{1},P_{2},\ldots,P_{n}\}$ be a set of projection operators satisfying $\sum_{i=1}^{n}P_{i}=\mathbf{1}\text{ and }P_{i}P_{j}=\delta_{ij}P_{i},$ where $\mathbf{1}$ denotes the identity operator. Suppose that an operator $O$ can be written as a linear combination of the projectors in $\mathcal{C}$, $O=\sum_{i=1}^{n}\alpha_{i}P_{i},$ with $\alpha_{i}\neq 0\,\forall\,i$. Then $O$ is invertible and its inverse $O^{-1}$ is given by
$O^{-1}=\sum_{i=1}^{n}\alpha_{i}^{-1}P_{i}.$
4  Solutions of the Problem
Since the wave operator in momentum space (1), where $\alpha(p),\beta(p),\gamma(p)\in\mathbb{C}$. Our main goal will be to invert it in the most general way possible. For this purpose, we shall begin a deeper study of the operators $\omega,\theta,S:\mathbb{C}^{3}\to\mathbb{C}^{3}$, defined by $(\omega x)^{\mu}=\omega^{\mu}_{\;\nu}x^{\nu}$, $(\theta x)^{\mu}=\theta^{\mu}_{\;\nu}x^{\nu}$ and $(Sx)^{\mu}=S^{\mu}_{\;\nu}x^{\nu}$, where $x=(x^{0},x^{1},x^{2})\in\mathbb{C}^{3}$, in order to obtain properties that will facilitate the obtaining of $\mathcal{O}^{-1}$.
4.1  Operators $\omega$, $\theta$ and $S$
These operators satisfy the following multiplication table.
Table 1: Products between $\omega$, $\theta$ and S.
$\omega$ $\theta$         S
$\omega$ $\omega$ 0         0
$\theta$ 0 $\theta$         S
S 0 S $p^{2}\theta$
Furthermore, it should be noted that $\omega$ and $\theta=I-\omega$ are orthogonal projectors that satisfy the following relations:
[…]
From theorems 3.1 and 3.2 we have $\mathrm{Im}\,\omega=\ker\theta$, $\ker\omega=\mathrm{Im}\,\theta$ and the decomposition $\mathbb{C}^{3}=\mathrm{Im}\,\omega+\mathrm{Im}\,\theta$.
Proposition 4.1.
The subspaces $\mathrm{Im}\,\omega$ and $W=\mathrm{Im}\,\theta=\ker\omega$ are invariant under $\omega,\theta$ and $S$. Let $S_{W}=S|_{W}:W\to W$ and $\mathbf{1}_{W}=\mathbf{1}_{W}:W\to W$ be the restrictions of $S$ and of the identity operator $\mathbf{1}$ to the subspace $W$. Then the operators $Z_{\pm}:W\to W$ defined by:
$Z_{\pm}=\frac{1}{2}\,\mathbf{1}_{W}\pm\frac{i}{2\sqrt{p^{2}}}\,S_{W}$ (9)
are projectors and therefore we have the decompositions
$W=\mathrm{Im}\,Z_{\pm}\oplus\ker Z_{\pm}.$
Moreover, such decomposition is non-trivial, that is,
$\dim(\mathrm{Im}\,Z_{\pm})=\dim(\ker Z_{\mp})=1,$
[…]
5  Calculation of the Propagators
First, let us begin by defining the following projection operators $\{P_{1}=\omega,P_{2}=Z_{+}\theta,P_{3}=Z_{-}\theta\}$. Or, we can write them explicitly as
$P_{1}=\omega,$
$P_{2}=\frac{1}{2}\,\theta+\frac{i}{2\sqrt{p^{2}}}\,S,$
$P_{3}=\frac{1}{2}\,\theta-\frac{i}{2\sqrt{p^{2}}}\,S.$
It should be noted that these operators satisfy the completeness relation
$\sum_{i}P_{i}=\mathbf{1}.$ (10)
And if we apply these operators to each other, we conclude that the orthogonality relationship is also satisfied
$P_{i}P_{j}=\delta_{ij}P_{i}.$ (11)
With such operators, one can construct the following multiplication table.
Table 2: Products between $P_{1}$, $P_{2}$ and $P_{3}$.
$P_{1}$ $P_{2}$ $P_{3}$
$P_{1}$ $P_{1}$ 0 0
$P_{2}$ 0 $P_{2}$ 0
$P_{3}$ 0 0 $P_{3}$
Now, writing the operators $\theta$ and $S$ in terms of the $P_{2}\text{ and }P_{3}$, we obtain
$\theta=P_{2}+P_{3},\quad S=-i\sqrt{p^{2}}(P_{2}-P_{3}).$ (12)
5.1  General Propagator
Substituting these relations into equation (1), we obtain
$\mathcal{O}=\alpha P_{1}+(\beta-i\gamma\sqrt{p^{2}})P_{2}+(\beta+i\gamma\sqrt{p^{2}})P_{3},$ (13)
Consequently, one can write the general propagator in terms of $P_{1},P_{2}$ and $P_{3}$,
$\mathcal{O}^{-1}=\frac{1}{\alpha}P_{1}+\frac{1}{\beta-i\gamma\sqrt{p^{2}}}P_{2}+\frac{1}{\beta+i\gamma\sqrt{p^{2}}}P_{3}.$ (14)
Thus, going back to $\omega,\theta$ and $S$, we obtain
$\mathcal{O}^{-1}=\frac{1}{\alpha}\,\omega+\frac{\beta}{\beta^{2}+\gamma^{2}p^{2}}\,\theta-\frac{\gamma}{\beta^{2}+\gamma^{2}p^{2}}\,S.$ (15)
5.2  The Maxwell-Lee-Wick-Chern-Simons Propagator
We start from the Lagrangian of Maxwell-Lee-Wick-Chern-Simons electrodynamics
$\mathcal{L}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{4m^{2}}F_{\mu\nu}\Box F^{\mu\nu}$
$\frac{1}{2\xi}(\partial^{\mu}A_{\mu})^{2}+\frac{\kappa}{2}\epsilon^{\mu\nu\rho}A_{\mu}\partial_{\nu}A_{\rho}-J_{\mu}A^{\mu},$ (16)
where $A^{\mu}$ is the vector potential, $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the electromagnetic field tensor and $J_{\mu}$ an external current. The gauge-fixing term, with the parameter $\xi$, must be included in order to make the calculation of the propagator possible. In fact, as will be shown, without this term the wave operator would not be invertible. In this way, using the coefficients of corresponding wave operator,
$\alpha=-\frac{p^{2}}{\xi},\quad\beta=\left(-p^{2}+\frac{p^{4}}{m^{2}}\right)\quad\text{and}\quad\gamma=i\kappa,$ (17)
we have the propagator for this theory
$\mathcal{O}^{-1}=-\frac{\xi}{p^{2}}P_{1}+\frac{1}{\left(-p^{2}+\frac{p^{4}}{m^{2}}\right)+\kappa\sqrt{p^{2}}}P_{2}$
$\frac{1}{\left(-p^{2}+\frac{p^{4}}{m^{2}}\right)-\kappa\sqrt{p^{2}}}P_{3}.$ (18)
Thus, at the base $\{\omega,\theta,S\}$, we have
$\mathcal{O}^{-1}_{\mu\nu}=-\frac{p^{2}\!\left(1-\frac{p^{2}}{m^{2}}\right)}{\,p^{2}\!\left[p^{2}\!\left(1-\frac{p^{2}}{m^{2}}\right)^{2}-\kappa^{2}\right]}\,\theta_{\mu\nu}$
$\frac{i\,\kappa}{\,p^{2}\!\left[p^{2}\!\left(1-\frac{p^{2}}{m^{2}}\right)^{2}-\kappa^{2}\right]}\,S_{\mu\nu}-\frac{\xi}{p^{2}}\omega_{\mu\nu}\,.$ (19)
5.3  The Maxwell-Deser-Jackiw Propagator
Another interesting model can be built up from the Maxwell system by adding the higher derivative Chern-Simons extension proposed by [deserjackiw]
$\mathcal{L}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+\frac{1}{2m}\,\varepsilon^{\mu\nu\lambda}(\Box A_{\mu})(\partial_{\nu}A_{\lambda})$
$\frac{1}{2\xi}(\partial_{\mu}A^{\mu})^{2}-J^{\mu}A_{\mu}.$ (20)
The corresponding wave operator has the following coefficients
$\alpha=-\frac{p^{2}}{\xi},\quad\beta=-p^{2}\quad\text{and}\quad\gamma=i\frac{p^{2}}{m}.$ (21)
Thus, the propagator of this theory is
$\mathcal{O}^{-1}=-\frac{\xi}{p^{2}}P_{1}-\frac{1}{p^{2}-\tfrac{p^{2}\sqrt{p^{2}}}{m}}P_{2}$
$\frac{1}{p^{2}+\tfrac{p^{2}\sqrt{p^{2}}}{m}}P_{3}.$ (22)
At the base $\{\omega,\theta,S\}$, with $\gamma=ip^{2}/m$, we have
$\mathcal{O}^{-1}_{\mu\nu}(p)=-\,\frac{1}{p^{2}\!\left(1-\frac{p^{2}}{m^{2}}\right)}\left(\theta_{\mu\nu}+\frac{i}{m}\,S_{\mu\nu}\right)-\frac{\xi}{p^{2}}\,\omega_{\mu\nu}.$ (23)
6  General analysis of causality and unitarity
In this section we present a systematic and model–independent analysis of causality and tree–level unitarity for the class of planar gauge theories whose wave operator in momentum space can be written as (1).
6.1   Analysis on the base $\{P_{1},P_{2},P_{3}\}$
Let $J_{\mu}(p)$ be an external conserved current, the saturated amplitude is then
$A(p)=J^{*\mu}(p)\,\mathcal{O}^{-1}_{\mu\nu}(p)\,J^{\nu}(p).$ (24)
Using the form (14), we obtain
$A(p)=\frac{\|J^{(+)}\|^{2}}{\beta(p)-i\gamma(p)\sqrt{p^{2}}}+\frac{\|J^{(-)}\|^{2}}{\beta(p)+i\gamma(p)\sqrt{p^{2}}},$ (25)
where the norms $\|J^{(\pm)}\|^{2}=(J^{(\pm)})^{*}\!\cdot J^{(\pm)}\geq 0$ are taken in the physical transverse subspace. Thus, we will call the coefficient of each term in (25) as follows
$F_{+}(p)=\frac{1}{\beta(p)-i\gamma(p)\sqrt{p^{2}}},$
$\qquad F_{-}(p)=\frac{1}{\beta(p)+i\gamma(p)\sqrt{p^{2}}}.$ (26)
6.1.1  Pole structure and general spectral analysis
The poles of the transverse propagator are determined by the zeros of the denominators in  (26). Multiplying the conjugate factors we find the common condition
$\bigl[\beta(p)-i\gamma(p)\sqrt{p^{2}}\bigr]\bigl[\beta(p)+i\gamma(p)\sqrt{p^{2}}\bigr]=$
$\beta(p)^{2}+\gamma(p)^{2}p^{2}=D(p)=0.$ (27)
Where $D(p)$ is given by
$D(p)\equiv\beta(p)^{2}+\gamma(p)^{2}p^{2}.$ (28)
Therefore all physical transverse poles are located at momentum $p^{\mu}=p_{*}^{\mu}$ satisfying
$D(p_{*})=0,$ (29)
i.e., at the solutions of
$\beta(p_{*})^{2}+\gamma(p_{*})^{2}\,p_{*}^{2}=0.$ (30)
For each fixed spatial momentum $\vec{p}$ these correspond to energies $p^{0}=E_{n}(\vec{p})$ given by the solutions of
$D(p^{0},\vec{p})=0.$ (31)
Real poles
If $E_{n}(\vec{p})$ is real, we obtain a propagating mode with dispersion relation determined implicitly by $D(E_{n},\vec{p})=0$. For such modes, causality at the microscopic level requires that the group velocity
$v_{g}=\left|\frac{\partial E_{n}}{\partial\vec{p}}\right|\leq 1,$ (32)
which can be translated into inequalities involving the derivatives of $D$ with respect to $p^{0}$ and $\vec{p}$ using
$\frac{\partial D}{\partial p^{0}}\Big|_{p_{*}}\,dE_{n}+\frac{\partial D}{\partial\vec{p}}\Big|_{p_{*}}\!\cdot d\vec{p}=0\quad\Rightarrow\quad\frac{\partial E_{n}}{\partial\vec{p}}=-\,\frac{\partial D/\partial\vec{p}}{\partial D/\partial p^{0}}\Bigg|_{p_{*}}.$ (33)
Complex poles
If some solutions $E_{n}(\vec{p})$ are complex, they occur in conjugate pairs by hermiticity of the underlying Lagrangian. Writing $E_{n}(\vec{p})=\Omega_{n}(\vec{p})-i\Gamma_{n}(\vec{p})$ with $\Gamma_{n}(\vec{p})>0$, the corresponding Fourier modes behave as $\exp[-i\Omega_{n}t]\,\exp[-\Gamma_{n}t]$ and can be interpreted this later on.
6.1.2  Residues and tree–level unitarity
To analyse unitarity we focus on the behaviour of the propagator near its poles. For a fixed $\vec{p}$, consider a simple pole at $p^{\mu}=p_{*}^{\mu}$ such that
$D(p_{*})=0,\qquad\left.\frac{\partial D}{\partial p^{0}}\right|_{p_{*}}\neq 0.$ (34)
Expanding $D$ around $p^{0}=E_{*}(\vec{p})$ we have
$D(p^{0},\vec{p})\simeq\left.\frac{\partial D}{\partial p^{0}}\right|_{p_{*}}\bigl(p^{0}-E_{*}(\vec{p})\bigr).$ (35)
Using (26), the scalar propagators near the pole behave as
$D_{\pm}(p^{0},\vec{p})\simeq\frac{\beta(p_{*})\pm i\gamma(p_{*})\sqrt{p_{*}^{2}}}{\left.\dfrac{\partial D}{\partial p^{0}}\right|_{p_{*}}}\,\frac{1}{p^{0}-E_{*}}.$ (36)
Therefore the residues of the saturated amplitudes are
$\text{Res}\,A_{\pm}\Big|_{p^{0}=E_{*}}=\|J^{(\pm)}\|^{2}\,\frac{\beta(p_{*})\pm i\gamma(p_{*})\sqrt{p_{*}^{2}}}{\left.\dfrac{\partial D}{\partial p^{0}}\right|_{p_{*}}},$ (37)
with $\|J^{(\pm)}\|^{2}\geq 0$. For real poles and real coefficients $\alpha(p)$, $\beta(p)$, $\gamma(p)$ on the real axis, the two residues are complex conjugates,
$\bigl(\text{Res}\,A_{+}\bigr)^{*}=\text{Res}\,A_{-}.$ (38)
Tree–level unitarity requires that the terms of (37) be non-negative.
6.1.3  Equivalent analysis in the base $\{\omega,\theta,S\}$
For completeness we show how the same conclusions follow directly from the form (15) of the propagator. Saturating with a conserved current $J_{\mu}$ and using $\omega J=0$ and $J=\theta J$, we obtain
$A(p)=J^{*}\!\cdot\mathcal{O}^{-1}\!\cdot J=\frac{\beta(p)}{D(p)}\,J^{*}\!\cdot J-\frac{\gamma(p)}{D(p)}\,J^{*}\!\cdot S\!\cdot J.$ (39)
In the transverse subspace we can again decompose $J$ into eigenvectors
$J=J^{(+)}+J^{(-)},\qquad SJ^{(\pm)}=\mp i\sqrt{p^{2}}\,J^{(\pm)}.$ (40)
It follows that
$J^{*}\!\cdot J=\|J^{(+)}\|^{2}+\|J^{(-)}\|^{2},$ (41)
and
$J^{*}\!\cdot S\!\cdot J=(J^{(+)})^{*}\!\cdot SJ^{(+)}+(J^{(-)})^{*}\!\cdot SJ^{(-)}=$
$i\sqrt{p^{2}}\,\|J^{(+)}\|^{2}+i\sqrt{p^{2}}\,\|J^{(-)}\|^{2}.$ (42)
Substituting into  (39) we obtain
$A(p)=\frac{1}{D(p)}\Bigl[\bigl(\beta(p)+i\gamma(p)\sqrt{p^{2}}\bigr)\,\|J^{(+)}\|^{2}\Bigr]$
$\frac{1}{D(p)}\Bigl[\bigl(\beta(p)-i\gamma(p)\sqrt{p^{2}}\bigr)\,\|J^{(-)}\|^{2}\Bigr],$ (43)
we recover exactly  (25). Thus the decomposition and the analysis of poles, residues, unitarity and causality are entirely equivalent in the $\{P_{1},P_{2},P_{3}\}$ and $\{\omega,\theta,S\}$ bases. The projector formalism therefore provides a compact and geometrically transparent framework for studying higher–derivative planar gauge models at the level of their propagators.
6.2   Analysis for the Maxwell–Lee–Wick–Chern–Simons model
Using (17), the physical poles of the transverse sector are therefore the solutions of
$p^{2}=0,\qquad p^{2}\!\left(1-\frac{p^{2}}{m^{2}}\right)^{2}-\kappa^{2}=0.$ (44)
The pole at $p^{2}=0\Rightarrow p^{0}=\pm\|\vec{p}\|$ is purely a gauge pole and disappears when the propagator is saturated with a conserved current, so that the physical structure is determined by the cubic term in $p^{2}$. For each fixed $\vec{p}$, defining $x=((p^{0})^{2}-\vec{p}^{\,2})/m^{2}$ and $\beta\equiv\kappa^{2}/m^{2}$,  (44) becomes
$x\left(1-x\right)^{2}-\beta=0.$ (45)
This cubic equation has three real roots for $0\leq\beta<4/27$, two degenerate real roots for $\beta=4/27$ and only one real root for $\beta>4/27$. In terms of the original parameters:
  • $\frac{\kappa^{2}}{m^{2}}<\frac{4}{27}$: three real massive modes (one of them with negative norm, as we will see below);
  • $\frac{\kappa^{2}}{m^{2}}=\frac{4}{27}$: critical case, with two degenerate modes;
  • $\frac{\kappa^{2}}{m^{2}}>\frac{4}{27}$: one real massive mode and one pair of complex-conjugate poles (Lee–Wick pair).
Each real root $x_{i}$ define $p_{i}^{2}=m^{2}x_{i}$. Therefore the dispersion relationship is $p^{0}=\pm\sqrt{\vec{p}^{\,2}+p_{i}^{2}}$. In other words, the physical modes are relativistic.
6.2.1  Saturated amplitude and residues
For a conserved current, the longitudinal sector ($P_{1}$) does not contribute. Using  (25) with the coefficients above, the saturated amplitude is
$A(p)=\frac{\|J^{(+)}\|^{2}}{-(p^{0})^{2}+\vec{p}^{\,2}+\frac{\big((p^{0})^{2}-\vec{p}^{\,2}\big)^{2}}{m^{2}}-\kappa\,\sqrt{(p^{0})^{2}-\vec{p}^{\,2}}}$
$\frac{\|J^{(-)}\|^{2}}{-(p^{0})^{2}+\vec{p}^{\,2}+\frac{\big((p^{0})^{2}-\vec{p}^{\,2}\big)^{2}}{m^{2}}+\kappa\,\sqrt{(p^{0})^{2}-\vec{p}^{\,2}}}.$ (46)
Defining the product of denominators $D(p^{0},\vec{p})=\left[(p^{0})^{2}-\vec{p}^{\,2}\right]G(p^{2})$, the residue of the saturated amplitude at $p^{0}=E^{*}$ is
$\text{Res}\,A\Big|_{p^{2}=\lambda}=-\,\frac{1-\lambda/m^{2}}{2G^{\prime}(\lambda)E^{*}}.$ (47)
Since,
$G^{\prime}(\lambda)=\left(1-\frac{\lambda}{m^{2}}\right)\left(1-\frac{3\lambda}{m^{2}}\right).$ (48)
How $E^{*}>0$, we see that the sign of the residue is controlled by
$\text{sgn}\,\bigl(\text{Res}\,A\bigr)=\text{sgn}\!\left[-\,\frac{1}{1-3\lambda/m^{2}}\right].$ (49)
Thus, for modes with $0<\lambda<m^{2}/3$ the residue is negative, characterizing a ghost mode (Lee–Wick partner), whereas modes with $\lambda>m^{2}/3$ have positive residue and correspond to physical unitary degrees of freedom. Adopting an appropriate contour prescription in the complex $p^{0}$ plane, these modes can be handled in such a way that the SS–matrix remains unitary and macroscopic causality is preserved, while possible microscopic violations of causality are confined to distances of order $1/\Gamma_{n}$ and are suppressed by the high mass scales characterizing the Lee–Wick partners.
6.2.2  Causality
For parameter values such that all poles $\lambda$ are real and positive ($\kappa^{2}/m^{2}<4/27$), the associated modes have relativistic dispersion relations and subluminal group velocities, preserving microscopic causality. When $\kappa^{2}/m^{2}>4/27$, two of the poles become complex conjugates; in this situation the coordinate-space propagator exhibits damped oscillations, corresponding to unstable resonances. Following the Lee–Wick contour prescription, one chooses the integration contour in $p^{0}$ so that such modes decay in the future, preserving the unitarity of the $S$-matrix and macroscopic causality, at the cost of small local violations of microcausality, suppressed by the Lee–Wick mass scale.
6.3   Analysis for the Maxwell–Deser–Jackiw model
Using (21), the poles are
$p^{4}=0\Rightarrow p^{2}=0,\qquad p^{2}=m^{2}.$ (50)
The pole at $p^{2}=0\Rightarrow p^{0}=\pm\|\vec{p}\|$ is purely a gauge pole and does not contribute to observables when the propagator is saturated with a conserved current. The physically relevant pole is therefore
[…]
which describes a single massive mode with mass $m>0$ and relativistic energy $E(\vec{p})$.
6.3.1  Saturated amplitude and residues
The saturated amplitude with a conserved current is obtained with $J=\theta J$ and $J^{*}\!\cdot\omega J=0$:
$A(p)=-\,\frac{1}{p^{2}\!\left(1-\dfrac{p^{2}}{m^{2}}\right)}\left(J^{*}\!\cdot J+\frac{i}{m}\,J^{*}\!\cdot S\cdot J\right).$ (52)
In the transverse subspace, we diagonalize $S$ in the basis
$SJ^{(\pm)}=\pm i\sqrt{p^{2}}\,J^{(\pm)},\qquad J=J^{(+)}+J^{(-)}.$ (53)
Hence,
$J^{*}\!\cdot J$ $=\|J^{(+)}\|^{2}+\|J^{(-)}\|^{2},$ (54)
$J^{*}\!\cdot S\cdot J$ $=+i\sqrt{p^{2}}\,\|J^{(+)}\|^{2}-i\sqrt{p^{2}}\,\|J^{(-)}\|^{2}.$
Substituting into (52),
$A(p^{0},\vec{p})=-\frac{1}{d(p^{0},\vec{p})}\Bigg[\|J^{(+)}\|^{2}\left(1-\frac{\sqrt{(p^{0})^{2}-\vec{p}^{\,2}}}{m}\right)\Bigg]$
[…]
At the physical pole we have
$A(p^{0},\vec{p})\Big|_{(p^{0})^{2}\simeq\vec{p}^{\,2}+m^{2}}\simeq-\,\frac{1}{d(p^{0},\vec{p})}\Big[\,0\cdot\|J^{(+)}\|^{2}+2\|J^{(-)}\|^{2}\Big].$ (56)
Where
$d(p^{0},\vec{p})=\big((p^{0})^{2}-\vec{p}^{\,2}\big)\left(1-\dfrac{(p^{0})^{2}-\vec{p}^{\,2}}{m^{2}}\right).$
Furthermore, one can also write
$A(p^{0},\vec{p})\Big|_{(p^{0})^{2}\simeq\vec{p}^{\,2}+m^{2}}\simeq\frac{\|J^{(-)}\|^{2}}{E^{*}(\vec{p})}\,\frac{1}{p^{0}-E^{*}(\vec{p})}$ (57)
This shows that only the negative term propagates, whereas the positive is projected out on the mass shell. The residue at $p^{o}=E^{*}(\vec{p})=\sqrt{\vec{p}^{\,2}+m^{2}}$ is
$\operatorname*{Res}_{\,p^{0}=E^{*}(\vec{p})}A(p^{0},\vec{p})=\frac{\|J^{(-)}\|^{2}}{E^{*}(\vec{p})}.$ (58)
How $E^{*}>0$ this demonstrates that the massive mode has positive norm and that the model is unitary at tree level.
6.3.2  Causality
Since the only physical pole is simple and real, with
$(p^{0})^{2}-\vec{p}^{\,2}=m^{2}\Rightarrow p^{0}=\pm E(\vec{p})\Rightarrow E(\vec{p})=\sqrt{\vec{p}^{\,2}+m^{2}}.$ (59)
For the positive–energy branch, the group velocity satisfies
$v_{g}=\left|\frac{\partial E(\vec{p})}{\partial|\vec{p}|}\right|=\frac{|\vec{p}|}{\sqrt{\vec{p}^{\,2}+m^{2}}}<1,$ (60)
in Lorentz units. So there is no superluminal propagation. Moreover, the absence of complex poles indicates that there are no Lee–Wick type modes; the operator $S_{\mu\nu}$ merely selects a single propagating therm without violating microcausality. Therefore, the Maxwell–Deser–Jackiw model describes a single massive mode that is both unitary and causal in $(2+1)$ dimensions.
7  Conclusions and Outlook
In this article we developed a systematic and fully general framework for studying causality and unitarity in planar gauge theories with higher-derivative and Chern–Simons terms. By expressing the momentum-space wave operator in the covariant $\{\omega,\theta,S\}$ basis with arbitrary momentum-dependent coefficients, we derived the most general form of the propagator and obtained explicit conditions on its pole structure, residues and analytic behavior. This produced a transparent set of criteria for identifying physical excitations, ghostlike modes and tachyons. The method was shown to reproduce and generalize the expected features of the Maxwell–Lee–Wick–Chern–Simons and Maxwell–Deser–Jackiw models, providing a unified perspective on their causal and unitary properties.
The causality and unitarity analysis can be equivalently formulated in the projector basis $(P_{1},P_{2},P_{3})$ or in the covariant basis $(\omega,\theta,S)$, both leading to the same physical results while emphasizing different aspects of the theory. The projector basis diagonalizes the wave operator in the transverse sector, making the spectral content explicit and allowing a direct identification of propagating modes, poles and unitarity properties, which is particularly useful for spectral analyses. The covariant basis, on the other hand, preserves manifest Lorentz covariance and provides a transparent interpretation of the operator structure by separating Maxwell, higher-derivative and Chern–Simons contributions, facilitating comparisons among different models. Although conceptually distinct, the two formulations are mathematically equivalent and preserve all pole positions and residues, thus offering complementary perspectives on the same underlying physics.
Several natural extensions of this work deserve further study. One promising direction is the application of the present framework to nonlocal or infinite-derivative generalizations of Maxwell–Chern–Simons theories, where the analytic structure of the propagator may display novel features such as nonpolynomial dispersion relations or softened UV behavior. Another way is the incorporation of interactions, either perturbatively or through resummation techniques, in order to determine how loop corrections modify the causal structure or shift the position of Lee–Wick poles. It would also be interesting to apply the method to gravitational models in $(2+1)$ dimensions, where higher-derivative and topological terms play a central role and where the spin-projector technique has proven particularly powerful.
Finally, the generalized analytic strategy developed here is well suited to examining whether causality and unitarity can coexist in more exotic settings, such as parity-violating dualities, Lorentz-violating effective actions, or finite-temperature backgrounds. We expect that the interplay between higher-derivative terms and topological masses in these regimes will continue to reveal rich structures, making planar gauge theories an ideal laboratory for probing fundamental aspects of quantum field theory.
Appendix A Supplementary Material
[…]
[Proof of Proposition 4.1] If $v\in\text{Im}\,\omega\implies v=\omega(w)\implies S(v)=S(\omega(w))=0$, since $S\omega=0$. Thus, we see that if $v\in\text{Im}\,\omega$, then $v\in\ker S$. In this way, the image of $\omega$ is invariant under $S$. It is easy to see that $\text{Im}\,\omega$ is also invariant under both $\omega$ and $\theta$. Indeed, note that if $v\in\text{Im}\,\omega$, then
$\omega(v)=v\in\text{Im}\,\omega,$
and
$\theta(v)=(I-\omega)(v)=I(v)-\omega(v)=v-\omega(v)=v-v=0.$
Now, setting $W=\text{Im}\,\theta=\ker\omega$, we show that $S(W)\subseteq W$. Let $w\in W$
$w=\theta(v)\implies S(w)=S(\theta(v))\implies S(w)=\theta(S(v))\in W,$
where we used the fact that $S$ and $\theta$ commute. From now on, we will only consider the restrictions of $S$ and $\theta$ to the subspace $W$ and we use the notation $S_{W}=S|_{W}:W\to W$ and $\theta_{W}=\theta|_{W}=\mathbf{1}_{W}$, so that
$S_{W}^{2}=-p^{2}\mathbf{1}_{W}.$
Defining $Z:W\to W$ as $Z=aS_{W}+b\mathbf{1}_{W}$, we want to find $a,b$ such that $Z$ is a projector, that is
$Z^{2}=(aS_{W}+b\mathbf{1}_{W})^{2}=Z,$
$a^{2}S_{W}^{2}+2abS_{W}+b^{2}\mathbf{1}_{W}=aS_{W}+b\mathbf{1}_{W},$
$\left(-a^{2}p^{2}+b^{2}\right)\mathbf{1}_{W}+2abS_{W}=aS_{W}+b\mathbf{1}_{W},$
thus,
$b=\frac{1}{2},\quad a=\frac{i}{2\sqrt{p^{2}}}.$
Therefore,
$Z=\frac{1}{2}\mathbf{1}_{W}+\frac{i}{2\sqrt{p^{2}}}S_{W},$
is a projector and, consequently, $W=\text{Im}\,Z\oplus\ker Z$. We can then write
$S_{W}=-\left(Z-\frac{1}{2}1_{W}\right)2i\sqrt{p^{2}}.$
If $w\in\text{Im}\,Z$, then
$S_{W}(w)=-\left(Z(w)-\frac{1}{2}1_{W}(w)\right)2i\sqrt{p^{2}}$
$\left(w-\frac{1}{2}w\right)2i\sqrt{p^{2}}=-i\sqrt{p^{2}}w=Sw.$
Similarly, if $w\in\ker Z$, then
$S_{W}(w)=i\sqrt{p^{2}}w=Sw.$
Therefore, we see that the vectors in $\text{Im}\,Z$ and $\ker Z$ are eigenvectors of $S$. Now, let us show that $\text{Im}\,Z$ and $\ker Z$ cannot be trivial subspaces! For this to be true, we must have $\dim(\text{Im}\,Z)=\dim(\ker Z)=1$. Suppose initially that one of them is trivial and we will arrive at a contradiction.
Case 1: $\text{Im}\,Z=W$. Thus, $\forall\,w\in W\implies\text{Im}\,Z=W$, i.e.
$S_{W}(w)=-i\sqrt{p^{2}}w\implies S_{W}=-i\sqrt{p^{2}}\mathbf{1}_{W}.$
Case 2: $\ker Z=W$. Thus, $\forall\,w\in W\implies\ker P_{W}=W$, i.e.
$S_{W}(w)=i\sqrt{p^{2}}w\implies S_{W}=i\sqrt{p^{2}}\mathbf{1}_{W}.$
In general, $S_{W}=a\mathbf{1}_{W}$. If $v=\omega(v)+\theta(v)$, then
$S(v)=S(\omega(v))+S(\theta(v))=S(\theta(v)),$
$S(v)=S_{W}(\theta(v)),$
$S(v)=a\mathbf{1}_{W}(\theta(v))\implies S(v)=a\theta(v),$
which gives us
$S=a\theta.$
Now let us make the following analysis,
$S^{\mu}_{\;\nu}=a\theta^{\mu}_{\;\nu},\quad S_{\mu\nu}=a\theta_{\mu\nu}.$
Since $\theta$ is symmetric, we can write
$S_{\mu\nu}=a\theta_{\nu\mu},$
but
$S_{\mu\nu}=S_{\nu\mu},\quad S_{\mu\nu}=-S_{\nu\mu},$
which implies $S_{\mu\nu}=0$, that is, $\theta=0$. But this is absurd, since $\theta\neq 0$. $\blacksquare$